# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ VI Acknowledgements
Appendix A Supplemental Material
A.1 Glossary of default RRE outputs
With the workflow of the graph-state-based fault-tolerant resource estimation method outlined in the main text, we summarize and reiterate the selected output parameters of RRE, providing new details in some cases. The parameters listed are RRE outputs to stdout and CSV files by default. We detail 48 logical architecture or hardware-relevant output parameters and explain how to calculate them. The delivered CSV file can include results for additional output parameters. Users can add custom classes of quantities for estimation to RRE using class templates from the resources.py module.
@@ A.1 Glossary of default RRE outputs
    $d$ denotes the surface code distance. All logical patches have a size of $2d^{2}$ and serve as foundational elements for all architectural components in the FTQC.
    We need to calculate the minimum viable $d$ precisely and iteratively according to Algorithm 1, which requires satisfying a logarithmic inequality in Step 6, also referred to in eq. 17.
@@
    The term $\tilde{n}_{\text{logical}}$ refers to the maximum number of logical nodes needed to process the complete graph state $\mathcal{G}$ according to the widgetization-based set $\tilde{S}_{\text{est}}$ at any given time. In other words, this represents the maximum quantum memory required. Each complete quantum memory register chain, shared among the modules on the same leg of the macro-architecture, contains $\tilde{n}_{\text{logical}}$ logical qubits. To prevent any quantum memory bottlenecks or delays, we allocate a fixed quantum memory and ancilla bus space of size $2\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$ to all modules and steps of the schedules. There may be more efficient methods for assigning the bilinear quantum register and ancilla bus spaces.
@@
  • Parameter 3: Number of $T$-distilleries (MSD factories) per module, $n_{\text{distilleries}}$
    The number of Litinski’s $T$-state distillation widgets or MSD factories per module is calculated as $n_{\text{distilleries}}=\lfloor\frac{l_{\text{edge}}-1}{\lfloor\frac{l_{\text{width}}}{2d^{2}}\rfloor}\rfloor n^{\text{col}}_{\text{distilleries}}$. The factories send purified magic states to $T$-transfer and ancillary buses necessary for the measurements on the logical qubits during the consumption stage. For more details, see section I.4.6.
@@
    The number of logical memory, or graph-state, or data qubits in the bilinear pattern in every module is equal to $\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$. These correspond to the blue tiles in fig. 6. In other words, every module on a leg of the macro-architecture contains this same number of logical memory qubits. Although the last module may require fewer active nodes for quantum operations, the same space is always allocated to this component.
@@
  • Parameter 6: Number of logical ancilla qubits in the ancilla bus per module
    The number of logical ancilla qubits in the ancilla bus section of each module is equal to $\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$. These correspond to the yellow tiles in fig. 6, which FTQC uses to facilitate logical measurements on the graph-state qubits. Each module on a leg of the macro-architecture contains the same number of logical ancilla qubits. Although the last module may need fewer active nodes for quantum operations, the same space is always allocated for this component.
  • Parameter 7: Number of physical ancilla qubits in the ancilla bus per module
    The number of physical ancilla qubits in the ancilla bus portion of every module is equal to $2d^{2}\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$.
@@
    The number of logical qubits in the $T$-transfer bus portion of every module is calculated as $l_{\text{transfer-bus}}=(l_{\text{edge}}-l_{\text{qbus}}-n^{\text{col}}_{\text{distilleries}}\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\rceil)l_{\text{edge}}+n^{\text{col}}_{\text{distilleries}}\lceil\frac{L_{\text{Length}}}{\sqrt{2}d}\rceil$. These represent the teal tiles in fig. 6, which FTQC utilizes to actively store and transfer purified $T$-states from MSD factories to the ancilla bus.
@@
  • Parameter 10: Number of logical qubits allocated to all $T$-distillation factories in each module
    The number of logical nodes assigned to all $n_{\text{distilleries}}$ factories in each module can be calculated as $n_{\text{distilleries}}\lceil L_{\text{length}}/(\sqrt{2}d)\rceil\lceil L_{\text{width}}/(\sqrt{2}d)\rceil$. These correspond to purple tiles in fig. 6.
  • Parameter 11: Number of physical qubits allocated to all $T$-distillation factories in each module
    The number of active physical qubits across all $n_{\text{distilleries}}$ factories per module can be computed as $n_{\text{distilleries}}Q$. After choosing a factory from the Distillation Widget Look-up table 1, $Q$ becomes the number of physical qubits the factory requires for MSD.
@@
    The total number of interconnected QPU cryo-modules on both legs of the macro-architecture can be computed as $2n^{\text{per-leg}}_{\text{modules}}$. The parameter $n^{\text{per-leg}}_{\text{modules}}$ is the number of modules per leg or the height of the ladder structure. See fig. 3 for details.
  • Parameter 13: Total number of interconnects in the macro-architecture
    The total number of width-$d$ coherent interconnects in the ladder macro-architecture can be computed as $n_{\text{inter-pipes}}(3n^{\text{per-leg}}_{\text{modules}}-2)$. FTQC uses these connections to perform inter-module graph preparation operations on the same-leg or teleporting output to input nodes on the opposite-leg modules. See fig. 3 for details.
@@
    The number of allocated physical qubits per module for all components can be computed as $2\lceil n_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil d^{2}+n_{\text{distilleries}}Q+2l_{\text{T-transfer}}d^{2}$.
@@
    The number of algorithmic logical qubits or width, $n_{\text{input}}$, of the complete input circuit, $U$. For the time-sliced widgetization we perform on the input algorithm, cf. fig. 7, all widgets have the same number of output and input nodes equal to $n_{\text{input}}$.
@@
    Extensive analytical and numerical studies (see, for example, [68, 69]) have established that a non-Clifford 1Q rotation gate can be systematically approximated as a chain of Clifford+$T$, a process known as gate synthesis. Here, the upper bound for $T$-length scales as $L_{\varepsilon}=\lceil c_{0}\log_{2}(1/\varepsilon)+c_{1}\rceil$. The diamond-norm precision, $\varepsilon$, indicates the rate at which FTQC utilizes the gate synthesis decomposition of arbitrary-angle rotations to perform $Rz$-basis measurements at the end, during each consumption step. We assume the use of the state-of-the-art decomposition approach of mixed-fallback [69] by default, which offers $c_{0}=0.57,c_{1}=8.83$, representing an improvement over the previously employed “GridSynth” method [68], which had $c_{0}=3.0,c_{1}=0$.
    RRE processes the arbitrary angle decompositions internally, which means the constructed graphs do not depend on $\varepsilon$-value and can be cached. The diamond-norm precision must be determined accurately and iteratively, following Algorithm 1, which naturally involves solving a logarithmic equation in Step 6.
    We argue that $\varepsilon$ should not be left for the user to set arbitrarily (although RRE allows users to impose a fixed value if needed). Extreme caution is required here because precisions are, indeed, equivalent to the fault-tolerant resources required: if $\varepsilon$ is chosen too large, the rate condition in Step 6 of Algorithm 1 is not met, while if it’s too small, the space-time resources will become wasteful and excessively large.
@@
    In the graph-state-based estimation method, our proxy for the total T-count, or $N^{\text{tot}}_{T}$, is the total number of T-basis measurements required after the arbitrary-angle $R_{z}$’s are synthesized at the measurement points of the consumption stage. Therefore, we can write T-count as $N^{\text{tot}}_{T}=N^{Rz}_{\text{init}}L_{\varepsilon}+N^{T}_{\text{init}}=N^{Rz}_{\text{init}}\lceil c_{0}\log_{2}(1/\varepsilon)+c_{1}\rceil+N^{T}_{\text{init}}$.
@@
    This parameter estimates the effective $T$-depth for the input logical circuit, $U$. This can be computed as $\lceil N^{T}_{tot}/\tilde{n}_{\text{logical}}\rceil$.
@@
    The number of arbitrary-rotation $Rz$-gates in the complete input algorithm, $U$, whether written explicitly or contained within other logical gates. All $Rz$ gates must undergo gate-synthesis decomposition [68, 69] to Clifford+$T$ at the end, during the consumption stage.
  • Parameter 20: Number of $T$,$T^{\dagger}$-gates in the input circuit, $N^{\text{init}}_{T}$
    The number of $T$ and $T^{\dagger}$ gates in the complete input algorithm, $U$, whether written explicitly or contained within other logical gates. We perform this count before performing gate-synthesis Clifford+$T$ decomposition of arbitrary-angle $R_{z}$-gates.
@@
    The number of explicit Clifford gates in the complete input algorithm, $U$.
  • Parameter 22: Number of graph state nodes, $N$
    $N$ represents the total number of logical nodes in the complete graph state, $G$, or the graph size. For widgetized circuits, we calculate a proxy for this quantity from subgraphs as $\tilde{N}=\sum_{i=0,...,n_{\text{widgets}}-1}N_{i}+\sum_{i=0,...,n_{\text{widgets}}-2}n^{\text{outputs}}_{i}\approx N$. Here, $n^{\text{outputs}}_{i}=n_{\text{input}}$ denotes the number of output nodes at step $i$ that must be added to the graph stitched from subgraphs to enable Bell pair teleportations.
@@
    The total number of sequential measurement steps in the consumption schedule over all widgets (repeated or not). In other words, the total number of subsets in the nested sets in the consumption schedule outputted by Cabaliser[71]. This quantity governs the quantum operation runtime for the graph consumption stage.
@@
    The total number of sequential measurement steps in the preparation schedule over all widgets (repeated or not). In other words, the total number of sub-sets in the nested sets in the preparation schedule outputted by Substrate Scheduler[44]. This quantity governs the quantum operation runtime for the graph preparation stage.
@@
    The total number of widgets (repeated or not), which is equivalent to the number of time steps in $U$ given the time-sliced decomposition presented in fig. 7 and section A.2.4.
@@
    $n^{*}_{\text{widgets}}$ is the number of distinct widgets in set $[U_{0},\dots,U_{n_{\text{widgets}}-1}]$ for the time-sliced widgetization of $U$. This means one can always write $n^{*}_{\text{widgets}}\leq n_{\text{widgets}}$. This is the number of widgets that FTQC needs to compile.
@@
    For our proposed superconducting fault-tolerant architecture, we assume it is feasible to engineer generic decoders utilizing modern GPU or FPGA hardware and methods such as MWPM or neural networks. $t^{\rm decoder}_{\rm tock}$ is the time required for the decoder to complete decoding $d$ QEC sweeps expressed in seconds. This can be approximated as $t^{\rm decoder}_{\rm tock}=dt_{\text{decoder}}$. We assume it is feasible to engineer generic GPU or FPGA-based decoders with a decoding tock or delay in the same order of magnitude as $t^{\rm decoder}_{\rm tock}$ for superconducting architecture. Here, we consider a characteristic decoder cycle time of $t_{\text{decoder}}=1~{}\mu$s.
@@
  • Parameter 29: Intra-module quantum tock for $T$-state distillations
    The time required for MSD factories to distill a $T$ state performing $d$ QEC sweeps sequentially. This can be computed as $8Ct$ for factories built from rotated surface code patches.
@@
    The number of logical qubits available per module allocated to different architectural components and quantum operations. This can be computed as $n_{\text{avail-logical}}=l_{\text{edge}}^{2}$.
  • Parameter 32: Total number of unallocated logical qubits
    Considering all modules and operations of the FTQC, the total number of unallocated logical qubits. This parameter can be computed as $2n_{\text{unalloc-logical}}n^{\text{per-leg}}_{\text{modules}}$, where $n_{\text{unalloc-logical}}=l_{\text{edge}}^{2}-2\lceil\frac{\tilde{n}_{\text{logical}}}{n^{\text{per-leg}}_{\text{modules}}}\rceil-l_{\text{T-transfer}}-n_{\text{distilleries}}\lceil\frac{l_{\text{length}}}{\sqrt{2}d}\rceil\lceil\frac{l_{\text{width}}}{\sqrt{2}d}\rceil$.
@@
    Considering all modules and operations of the FTQC, the total number of allocated logical qubits to different architectural components. This parameter can be computed as $2n_{\text{alloc-logical}}n^{\text{per-leg}}_{\text{modules}}$, where $n_{\text{alloc-logical}}=2\lceil\frac{\tilde{n}_{\text{logical}}}{n^{\text{per-leg}}_{\text{modules}}}\rceil+l_{\text{T-transfer}}+n_{\text{distilleries}}\lceil\frac{l_{\text{length}}}{\sqrt{2}d}\rceil\lceil\frac{l_{\text{width}}}{\sqrt{2}d}\rceil$.
@@
    For our superconducting architecture, we assume generic reference classical decoders equipped with state-of-the-art GPU or FPGA cores perform decoding cycles. These receive the syndrome and output possible corrective operations, which can be, for example, performed at the same time of measurements at the ends of consumption steps. Quantum operations do not need to wait for the generic decoders we have assumed to finish their cycles to perform corrective measurements. One can always consider a staggered type of architecture, where classical and quantum units work together at each step. Moreover, since decoder tocks are typically much longer than quantum tocks, the best strategy is to let as many concurrent decoding cores complete their cycles as possible at each consumption step. In this way, the decoding process only adds an overall delay to the consumption stage runtime on top of purely quantum operations. There would be a maximum number of cores that can run concurrently at any consumption step.
    We define $\text{cores}_{\text{consump}}^{\text{decoding}}$ as the maximum number of concurrent decoding cores running at any consumption step. Hence, consumption steps often dominate intramodule runtime, we set $\text{cores}_{\text{consump}}^{\text{decoding}}$ as the available number of concurrent cores for all decoding steps, including graph preparation, teleportation, consumption, or $T$ injection operations. This is calculated as $\text{cores}_{\text{consump}}^{\text{decoding}}=\lceil\frac{t_{\text{tock}}^{\text{decoder}}}{t^{\text{intra}}_{\text{tock}}}\rceil$.
@@
  • Parameter 40: Power dissipation at 4K stage, $\text{POW}_{\text{4K}}$
    Power dissipation for the 4-Kelvin stages of all cryo-modules in Watts.
  • Parameter 41: Total graph consumption time, $t^{\rm tot}_{\rm consump}$
@@
    The total time to consume subgraphs across all consumption schedule steps is expressed in seconds. $t^{\rm tot}_{\rm consump}$ includes the delays required to distill additional $T$-states or prepare the subgraphs, while excluding handover runtime and decoding delays. This is calculated as $t^{\rm tot}_{\rm consump}=8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}$, as detailed in section I.4.6.
@@
  • Parameter 42: Total inter-modular graph handover time, $t^{\text{tot}}_{\text{handover-inter}}$
@@
    The total inter-modular time for Bell-state teleportations to handover (stitch) subgraphs for all scheduling steps. This is the time to teleport the output nodes on a current subgraph on the quantum register on one leg to the input nodes of the subgraph on the other leg for all steps. This is calculated as $t^{\text{tot}}_{\text{handover-inter}}=\sum_{i=0,\dots,n_{\text{widgets}}-2}t^{i}_{\text{handover-inter}}$ and was detailed in section I.4.6.
  • Parameter 43: Overall $T$-distillation delay, $t^{\text{distill-delay}}_{tot}$
@@
    $t^{\text{distill-delay}}_{tot}$ is the overall delay included in the total consumption time, $t^{\rm tot}_{\rm consump}$, to distill additional $T$-states in the MSD factories in the same modules to complete consumption-stage measurements. See section I.4.6 for details.
  • Parameter 44: Overall graph preparation delay, $t^{\text{prep-delay}}_{tot}$
@@
    $t^{\text{prep-delay}}_{tot}$ is the overall delay included in the total consumption time, $t^{\rm tot}_{\rm consump}$, to create the next subgraph in the schedule. The FTQC pipeline cannot proceed from the current consumption step until the next subgraph is prepared in the next set of modules on the other leg of the ladder. See section I.4.6 for details.
  • Parameter 45: Overall decoding delay, $t^{\text{decode-delay}}_{tot}$

### v3

@@ VII Acknowledgments
VIII DATA AVAILABILITY
The data that supports the findings of this article cannot be made publicly available. The data are available upon reasonable request from the authors. Codes, scripts, and sample configuration files required to reproduce similar results are accessible via the RRE software [77].
Appendix A Glossary of default RRE outputs
Building on the workflow of the graph-state-based fault-tolerant resource estimation method outlined in the main text, we summarize and reiterate the selected output parameters of RRE, providing additional details where needed. We detail $49$ logical architecture parameters or hardware-relevant outputs, which RRE writes to stdout and CSV files by default, and explain how to calculate them. The delivered CSV file can include results for additional output parameters. Users can add custom quantity classes for estimation in RRE using class templates from the resources.py module.
@@ Appendix A Glossary of default RRE outputs
    $d$ denotes the surface code distance of logical patches. For the surface code, $\left\lfloor(d-1)/2\right\rfloor$ errors can be corrected before a logical operation fails. All non-distillation logical patches have size $2d^{2}$ and serve as foundational elements for architectural components in the FTQC ($T$-factories are parameterized by their own code-distance sets).
    We need to calculate the minimum viable $d$ precisely and iteratively according to Algorithm 1, which requires satisfying a logarithmic inequality in Step 6, also referred to in eq.˜17.
@@
    The term $\tilde{n}_{\text{logical}}$ refers to the maximum number of logical nodes needed to process the complete graph state $\mathcal{G}$ according to the widgetization-based set $\tilde{S}_{\text{est}}$ at any given time. In other words, this represents the maximum quantum memory required. Each complete quantum memory register chain, shared among the modules on the same leg of the macro-architecture, contains $\tilde{n}_{\text{logical}}$ logical qubits. To prevent any quantum memory bottlenecks or delays, we allocate a fixed quantum memory and auxiliary bus space of size $2\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$ to all modules and steps of the schedules. There may be more efficient methods for assigning the bilinear quantum register and auxiliary bus spaces.
@@
  • Parameter 3: Number of $T$-factories per module, $n_{\text{T-factories}}$
    The number of Litinski-style $T$-factories per module is calculated as $n_{\text{T-factories}}=\left\lfloor\frac{l_{\text{edge}}-1}{\left\lfloor\frac{l_{\text{width}}}{2d^{2}}\right\rfloor}\right\rfloor n^{\text{col}}_{\text{T-factories}}$. The factories send purified magic states to the $T$-transfer and auxiliary buses required for measurements on the logical qubits during the consumption stage. For more details, see section˜II.4.7.
@@
    The number of logical memory, or graph-state, or data qubits in the bilinear pattern in every module is equal to $\left\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\right\rceil$. These correspond to the blue tiles in fig.˜1. In other words, every module on a leg of the macro-architecture contains the same number of logical memory qubits. Although the last module may require fewer active nodes for quantum operations, the same space is always allocated to this component.
@@
  • Parameter 6: Number of logical auxiliary qubits in the auxiliary bus per module
    The number of logical auxiliary qubits in the auxiliary bus section of each module is equal to $\left\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\right\rceil$. These correspond to the yellow tiles in fig.˜1, which FTQC uses to facilitate logical measurements on the logical graph-state qubits. Each module on a leg of the macro-architecture contains the same number of logical auxiliary qubits. Although the last module may need fewer active nodes for quantum operations, the same space is always allocated for this component.
  • Parameter 7: Number of physical qubits in the auxiliary bus per module
    The number of physical qubits in the auxiliary bus portion of every module is equal to $2d^{2}\left\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\right\rceil$.
@@
    The number of logical qubits in the $T$-transfer bus portion of every module is calculated as $l_{\text{transfer-bus}}=(l_{\text{edge}}-l_{\text{qbus}}-n^{\text{col}}_{\text{T-factories}}\left\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\right\rceil)l_{\text{edge}}+n^{\text{col}}_{\text{T-factories}}\left\lceil\frac{L_{\text{Length}}}{\sqrt{2}d}\right\rceil$. These correspond to the teal tiles in fig.˜1, which FTQC uses to actively store and transfer purified $T$-states from MSD factories to the auxiliary bus.
@@
  • Parameter 10: Number of logical qubits allocated to all $T$-factories in each module
    The number of logical nodes assigned to all $n_{\text{T-factories}}$ $T$-factories in each module can be calculated as $n_{\text{T-factories}}\left\lceil L_{\text{length}}/(\sqrt{2}d)\rceil\lceil L_{\text{width}}/(\sqrt{2}d)\right\rceil$. These correspond to purple tiles in fig.˜1.
  • Parameter 11: Number of physical qubits allocated to all $T$-factories in each module
    The number of active physical qubits across all $n_{\text{T-factories}}$ factories per module can be computed as $n_{\text{T-factories}}Q$. After choosing a factory from the $T$-factory Look-up Table 1, $Q$ becomes the number of physical qubits the factory requires for magic state distillation.
@@
    The total number of interconnected QPU cryo-modules on both legs of the macro-architecture can be computed as $2n^{\text{per-leg}}_{\text{modules}}$. The parameter $n^{\text{per-leg}}_{\text{modules}}$ is the number of modules per leg or the height of the ladder structure. See fig.˜3 for details.
  • Parameter 13: Total number of coherent interconnects in the macro-architecture
    The total number of width-$d$ coherent interconnects in the ladder macro-architecture can be computed as $n_{\text{inter-pipes}}\left(3n^{\text{per-leg}}_{\text{modules}}-2\right)$. FTQC uses these connections to perform inter-module graph preparation operations on the same-leg or teleporting output to input nodes on the opposite-leg modules. See fig.˜3 for details.
@@
    The number of allocated physical qubits per module for all components can be computed as $2\left\lceil n_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\right\rceil d^{2}+n_{\text{T-factories}}Q+2l_{\text{transfer-bus}}d^{2}$.
@@
    The number of algorithmic logical qubits or width, $n_{\text{input}}$, of the complete input circuit, $U$. For the time-sliced widgetization we perform on the input algorithm (cf. fig.˜6), all widgets have the same number of input and output nodes, $n_{\text{input}}$.
@@
    Extensive analytical and numerical studies (see, for example, [75, 32]) have established that a non-Clifford 1Q rotation gate can be systematically approximated as a chain of Clifford+$T$, a process known as gate synthesis. Here, the upper bound for $T$-length scales as $L_{\varepsilon}=\left\lceil c_{0}\log_{2}(1/\varepsilon)+c_{1}\right\rceil$. The diamond-norm precision, $\varepsilon$, indicates the rate at which FTQC utilizes the gate synthesis decomposition of arbitrary-angle rotations to perform $Rz$-basis measurements at the end, during each consumption step. We assume the use of the state-of-the-art decomposition approach of mixed-fallback [32] by default, which offers $c_{0}=0.57,c_{1}=8.83$, representing an improvement over the previously employed “GridSynth” method [75], which had $c_{0}=3.0,c_{1}=0$.
    RRE internally processes non-Clifford angle decompositions, meaning the constructed graphs do not depend on the $\varepsilon$-value and can be cached. The diamond-norm precision must be determined accurately and iteratively, following Algorithm 1, which naturally involves solving a logarithmic equation in Step 6.
    We argue that $\varepsilon$ should not be left for the user to set arbitrarily (although RRE allows users to impose a fixed value if needed). Extreme caution is required here because precisions are, indeed, equivalent to the fault-tolerant resources required: if $\varepsilon$ is chosen too large, the rate condition in Step 6 of the algorithm 1 is not met, while if it’s too small, the space-time resources will become wasteful and excessively large.
@@
    In the graph-state-based estimation method, our proxy for the total T-count, or $N^{\text{tot}}_{T}$, is the total number of $T$-basis measurements required after the arbitrary-angle $R_{z}$’s are synthesized at the measurement points of the consumption stage. Therefore, we can write T-count as $N^{\text{tot}}_{T}=n^{Rz}_{\text{init}}L_{\varepsilon}+n^{T}_{\text{init}}=n^{Rz}_{\text{init}}\left\lceil c_{0}\log_{2}(1/\varepsilon)+c_{1}\right\rceil+n^{T}_{\text{init}}$.
@@
    This parameter estimates the effective $T$-depth for the input logical circuit, $U$. This can be computed as $\left\lceil n^{T}_{tot}/\tilde{n}_{\text{logical}}\right\rceil$.
@@
    We denote the number of non-Clifford-rotation RzRz-gates in the complete input algorithm, $U$, whether written explicitly or contained within other logical gates, as $n^{\text{init}}_{Rz}$. All non-Clifford $Rz$ gates must undergo gate-synthesis decomposition [75, 32] to Clifford+$T$ at the end, during the consumption stage.
  • Parameter 20: Number of $TT^{\dagger}-$gates in the input circuit, $n^{\text{init}}_{T}$
    We denote the number of $T$ and $T^{\dagger}$ gates in the complete input algorithm, $U$, whether written explicitly or contained within other logical gates, as $n^{\text{init}}_{T}$. We perform this count before performing gate-synthesis Clifford+$T$ decomposition of non-Clifford $R_{z}$-gates.
@@
    The number of explicit Clifford gates in the complete input algorithm, $U$, which is denoted as $n^{\text{init}}_{\text{Clifford}}$.
  • Parameter 22: Number of graph state nodes, NN
    NN represents the total number of logical nodes in the complete graph state, $\mathcal{G}$, or the graph size. For widgetized circuits, we calculate a proxy for this quantity from subgraphs as $\tilde{N}=\sum_{i=0,...,n_{\text{widgets}}-1}N_{i}+\sum_{i=0,...,n_{\text{widgets}}-2}n^{\text{outputs}}_{i}\approx N$. Here, $n^{\text{outputs}}_{i}=n_{\text{input}}$ denotes the number of output nodes at step $i$ that must be added to the graph stitched from subgraphs to enable Bell pair teleportations.
@@
    The total number of sequential measurement steps in the consumption schedule over all widgets (repeated or not). In other words, the total number of subsets in the nested sets in the consumption schedule outputted by Cabaliser[72]. This quantity governs the runtime of the quantum operation for the graph consumption stage.
@@
    The total number of sequential measurement steps in the preparation schedule over all widgets (repeated or not). In other words, the total number of sub-sets in the nested sets in the preparation schedule outputted by Substrate Scheduler[41]. This quantity governs the runtime of the quantum operation for the graph preparation stage.
@@
    The total number of widgets (repeated or not), which is equivalent to the number of time steps in $U$ given the time-sliced decomposition presented in fig.˜6 and section˜B.0.4.
@@
    $n^{*}_{\text{widgets}}$ is the number of distinct widgets in set $U_{0},\dots,U_{n_{\text{widgets}}-1}]$ for the time-sliced widgetization of $U$. This means one can always write $n^{*}_{\text{widgets}}\leq n_{\text{widgets}}$. This is the number of widgets that FTQC needs to compile.
@@
[…]
@@
  • Parameter 29: Intra-module quantum tock for $T$-factories
    The time required for $T$-factories to distill a $T$ state performing $d$ QEC sweeps sequentially. This can be computed as $8Ct$ for $T$-factories built from rotated surface code patches.
@@
    The number of logical qubits available per module, which is allocated to different architectural components and quantum operations. This can be computed as $n_{\text{avail-logical}}=l_{\text{edge}}^{2}$.
  • Parameter 32: Total number of unallocated logical qubits, $n_{\text{unalloc-logical}}$
    Considering all modules and operations of the FTQC, the total number of unallocated logical qubits. This parameter can be computed as $2n_{\text{unalloc-logical}}n^{\text{per-leg}}_{\text{modules}}$, where $n_{\text{unalloc-logical}}=l_{\text{edge}}^{2}-2\left\lceil\frac{\tilde{n}_{\text{logical}}}{n^{\text{per-leg}}_{\text{modules}}}\right\rceil-l_{\text{transfer-bus}}-n_{\text{T-factories}}\left\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\right\rceil\left\lceil\frac{L_{\text{width}}}{\sqrt{2}d}\right\rceil$.
@@
    Considering all modules and operations of the FTQC, the total number of allocated logical qubits to different architectural components. This parameter can be computed as $2n_{\text{alloc-logical}}n^{\text{per-leg}}_{\text{modules}}$, where $n_{\text{alloc-logical}}=2\left\lceil\frac{\tilde{n}_{\text{logical}}}{n^{\text{per-leg}}_{\text{modules}}}\right\rceil+l_{\text{transfer-bus}}+n_{\text{T-factories}}\left\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\right\rceil\left\lceil\frac{L_{\text{width}}}{\sqrt{2}d}\right\rceil$.
@@
    In our superconducting architecture, we assume that generic reference classical decoders, equipped with state-of-the-art GPU or FPGA cores, perform decoding cycles. These receive the syndrome and output possible corrective actions, which can, for example, be performed at the same time as measurements at the ends of consumption steps.
    Quantum operations do not need to wait for the generic decoders we have assumed to finish their cycles to perform corrective measurements. One can always consider a staggered type of architecture, where classical and quantum units work together at each step. Moreover, since decoder tocks are typically much longer than quantum tocks, the best strategy is to let as many concurrent decoding cores complete their cycles as possible at each consumption step. In this way, the decoding process only adds an overall delay to the consumption stage runtime on top of purely quantum operations. There would be a maximum number of cores that can run concurrently at any consumption step.
    We define $\text{cores}_{\text{consump}}^{\text{decoding}}$ as the maximum number of concurrent decoding cores running at any consumption step. Hence, consumption steps often dominate intramodule runtime, so we set $\text{cores}_{\text{consump}}^{\text{decoding}}$ as the available number of concurrent cores for all decoding steps, including graph preparation, teleportation, consumption, or $T$ injection operations. This is calculated as $\text{cores}_{\text{consump}}^{\text{decoding}}=\left\lceil\frac{t_{\text{tock}}^{\text{decoder}}}{t^{\text{intra}}_{\text{tock}}}\right\rceil$.
@@
  • Parameter 40: Power dissipation at 4K stage, $P_{\text{Diss},\text{4K}}$
    Power dissipation for the $4$-Kelvin stages of all (cryo-)modules in Watts.
  • Parameter 41: Power dissipation at 20mK stage, $P_{\text{Diss},\text{20mK}}$
    Power dissipation for the $20$-milliKelvin stages of all (cryo-)modules in Watts.
  • Parameter 42: Total graph consumption time, $t^{\rm tot}_{\rm consump}$
@@
    The total time to consume subgraphs across all consumption schedule steps is expressed in seconds. $t^{\rm tot}_{\rm consump}$ includes the delays required to distill additional $T$-states or prepare the subgraphs, while excluding handover runtime and decoding delays. This is calculated as $t^{\rm tot}_{\rm consump}=8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}$, as detailed in section˜II.4.7.
@@
  • Parameter 43: Total inter-modular graph handover time, $t^{\text{tot}}_{\text{handover-inter}}$
@@
    The total inter-modular time for Bell-state teleportations to hand over (stitch) subgraphs across all scheduling steps. This is the time to teleport the output nodes on the current subgraph on the quantum register on one leg to the input nodes of the subgraph on the other leg for all steps. This is calculated as $t^{\text{tot}}_{\text{handover-inter}}=\sum_{i=0,\dots,n_{\text{widgets}}-2}t^{i}_{\text{handover-inter}}$ and was detailed in section˜II.4.7.
  • Parameter 44: Overall $T$-distillation delay, $t^{\text{distill-delay}}_{tot}$
@@
    $t^{\text{distill-delay}}_{tot}$ is the overall delay included in the total consumption time, $t^{\rm tot}_{\rm consump}$, to distill additional $T$-states in the $T$-factories in the same modules to complete consumption-stage measurements. See section˜II.4.7 for details.
  • Parameter 45: Overall graph preparation delay, $t^{\text{prep-delay}}_{tot}$
@@
    $t^{\text{prep-delay}}_{tot}$ is the overall delay included in the total consumption time, $t^{\rm tot}_{\rm consump}$, to prepare the next subgraph in the schedule. The FTQC pipeline cannot proceed from the current consumption step until the next subgraph is prepared in the next set of modules on the other leg of the ladder. See section˜II.4.7 for details.
  • Parameter 46: Overall decoding delay, $t^{\text{decode-delay}}_{tot}$